a)     Draw a production function that exhibits diminishing marginal product of labor. Define total cost, average total cost, and marginal cost. How are they related?


1
Expert's answer
2021-07-06T11:55:11-0400

Production function with diminishing marginal product of labour



Total cost is the overall sum of production cost comprising of both explicit and implicit cost spent to produce a certain level of output. 


Average total cost is the total expenditure incurred over the total quantity produced. 


Marginal cost is the change in total cost incurred  when output increases by a single unit.


How they are related

AC = "\frac{TC}{Q}" ; TC = AC× Q

"MC =\frac{ \Delta TC}{\Delta Q}"